OpenStudy (anonymous):

how do you decompose 3x^2+7x+1/x(x+1)^2 into partial fractions?

OpenStudy (anonymous):

\[\frac{3x^2+7x+1}{x(x+1)^2}\] \[\frac{a}{x}+\frac{b}{x+1}+\frac{c}{(x+1)^2}\] is a start

OpenStudy (anonymous):

then maybe \[a(x+1)^2+bx(x+1)+cx=3x^2+7x+1\] you get \(a\) immediately by replacing \(x\) by 0

OpenStudy (anonymous):

in fact you get \(a=1\) since the left side is \(a\) and the right side is \(1\)

OpenStudy (anonymous):

now replace \(x\) by \(-1\) and get \[-c=-3\] so \(c=3\)

OpenStudy (anonymous):

I can't wait to learn this properly.

OpenStudy (anonymous):

@eliassaab has a snap way to do this that i cannot remember, but now that we have \(a=1\) and \(c=3\) we can find \(b\) any number of ways. we see that the first term gives \(x^2\) and the second will give \(bx^2\) which together must be \(3x^2\) making \(b=2\)

OpenStudy (anonymous):

i am not sure what "properly" means but there are many ways, on involving derivatives and the other limits, but i forget them. this is the easiest way i know how you can also multiply out and equate like coefficients, but usually that is a long drag perhaps not in this case

OpenStudy (anonymous):

i dont understand how you went from the first part to a(x+1)^2+bx(x+1)+cx=3x^2+7x+1

OpenStudy (anonymous):

add the fractions and that will be your numerator

OpenStudy (anonymous):

@satellite73 : I mean't I can't wait till I learn this in Class. So far I learnt Partial fractions through online videos.